# MetMamba: Regional Weather Forecasting with Spatial-Temporal Mamba Model

### v1

@@ 1 Introduction
Trained on European Centre for Medium-range Weather Forecast (ECMWF)’s reanalysis product ECMWF Reanalysis v5 (ERA5) dataset, deep learning weather forecasting models (DLWP) have shown promising performance, with FourCastNet[5] being the first data-driven model that directly competes with ECMWF’s Integrate Forecasting System (IFS), followed by a series of other models[6][7][8][9] that outperforms IFS in various metrics [10], these models demonstrates excellent ability in capturing global weather trends, with a fraction of the computation requirement.
Though a coarser grid global forecast is crucial for gaining meaningful insights into longer horizon outlooks, extreme weather events often come in the form of convection, which are sub-grid dynamics not reflected in coarser grid data, even with advancements in fine grid global forecast, parameterization of such events also requires localized tunning[11], making limited area model (LAM) pivotal in many applications. We conduct this research within the LAM context, partly due to its greatly reduced compute, but also aim to explore the feasibility of deep learning for Limited Area Modeling (LAM).
@@ 2.2 Transformers
Transformers, known for its ability to associate information in various parts of the input data via the attention mechanism[18], are initially designed for natural language processing tasks, have quickly gain wide adoption in computer vision with Vision Transformers[19]. Its quadratic complexity (1) however, poses as a significant hurdle in its application, particularly in vision. Different approaches to alleviate this issue have been proposed, one most popular among them is Swin-Transformer[20], which uses a switching window mechanism to reduce the number of required tokens to achieve a global receptive field. Sub-quadratic approximations of attention like Linformer[21] are also a viable surrogate. In the DLWP space, where the compute efficiency is of a lesser concern, the use of transformers is ubiquitous[7][8][22][9][23], favoured by its modeling power. But as noted by [23], the patching step which are used to reduce the number of tokens also leads to degradation in performance.
@@ 2.3 Mamba Model
Mamba[12] model is a class of State Space Models (SSMs), which is originally designed to process continuous signals, bearing close similarity in formulations to RNNs. (Eq.(2)) Its primary contribution, aside from parallelization and hardware level optimization, is enabling SSMs to process inputs in a time dependant manner, making it a viable alternative to attention [24] Modifying it for vision use cases involves transferring its 1D-Scan pattern to the 2d space[13][25], which achieved better performance than other leading vision transformer models.
@@ 3.1 Dataset
For the lateral boundary condition, we chose the latest iteration of FourCastNet (SFNO), distributed with Earth2Studio[29]. Among models with wide adoption and benchmarks, it is the only one with a continuous 6 hour rollout scheme and specific humidity as its vapor related variable. The LBCs are produced by initializing the model with ERA5 global data from 2018 onward, making a relatively smaller (2018-2021) for LBCs adaptation training, avoiding the years where the global model was trained on. See Appendix A for details regarding the dataset used.
Figure 1: (a) general model architecture: the DLWP-LAM takes primary inputs: two initial conditions $IC_{0}$ at initialization time $t_{0}$ and $IC_{-1}$ at previous (6 hours before) initialization time $t_{0-6h}$, and auxiliary inputs: constants from Digital Elevation Model (DEM) and elapsed time, they are processed by a DEM specific embedder, a conditioning module or by simple concatenation. The input tensor is then processed by $N$ blocks. The output is decoded with a simple linear layer and pixel shuffle operation. (b) DEM Embedder: an embedder that process different information from the DEM accordingly. (c) AdaLN Block: a block that integrates elapsed year time via the use of adaLN, this will regulate all blocks in the process step.
@@ 3.2 Models
This is a Swin-Transformer-V2 block used in Swin-Transformer-V2[30] and also by FuXi[8], with no modifications, we refer to their paper for implementation details.
@@
We design a dual branch model that mixes latent information between Fourier Neural Operator and SwinV2 branches using a FFT cross attention mechanism [31], achieving better results than using either of the single branch models, while alleviating the limitation of AFNO in the limited area setting.
@@
We adopt the cross-scan and cross-merge methods used by VMamba[13], and extend these methods to the 3 dimensional space, to better integrate information from spatial and temporal dimensions, similarly, a 3D depth-wise convolution module is also used. We delineate the design choices of Mamba-3D in Appendix B.2.3
@@ 3.6 Evaluation
We evaluate all LAM models, FourCastNet(SFNO) and relevant ablation studies with initial conditions from test year 2021. In the total 1439 initial conditions, only 1418 of them are used, with the first 2 and last 20 initial conditions reserved for model initialization and forecast truncation. The models are run and evaluated for 20 steps. Only Root Mean Squared Error (RMSE) is provided as a performance metric for simplicity. We note that this metric is calculated following the convention by WeatherBench[34], with averaging as the last step of the calculation, latitude weighting scheme is not used.
@@ 3.7 Environment
These models are trained on two Nvidia A100-40GB cards utilizing PyTorch’s Fully Sharded Distributed Parallel, gradient checkpointing and Automatic Mixed Precision.
@@ 5 Discussion
The Mamba block, when used as a drop-in replacement in our existing LAMs offers a notable performance uplift. This can be attributed to its ability to both generate global receptive field and associate nearby tokens more efficiently than convolution networks. A reduced direct association between disjointed and distant tokens than vision transformers, which we suspect is less common in the field of meteorology, could also be a contributing factor. Its mathematical formulation being more suited to process continuous signals is another valid hypothesis. We also note that this scan pattern is natively transferable and more suitable to a global geometry.
@@
: We purpose Previous models have used simple linear layers to process elevation, soil, land-sea mask and other time related
@@ Appendix A Data Details
Digital Elevation Model data is sampled from the Weather Research and Forecast (WRF) model, containing land-sea mask, soil type, snow albedo etc. Positional information is added by calculating the sine and cosine normalized latitude and longitude. (See Eq.(A) and Eq.(6)) Time of day is added as a group of 2 dimensional sine and cosine normalized local time. Time of year is added as a group of 1 dimensional sine and cosine normalized current UTC time (in seconds) divided by the total seconds per year.
@@ Appendix B Model Details
We delineate the design and implementation details of the our MetMamba model in this section. The main difference to two other types of DLWP-LAMs is MetMamba’s ability to process spatial-temporal inputs natively, without the need for an additional 3-D convolution encoder. Aside from sharing the same structure, auxiliary information encoder and decoder, all DLWP-LAMs share the same embedding size and similar parameter count. The SwinV2 based LAM serves as a baseline, the AFNO-Swin based LAM serves as a incremental update with relatively higher efficiency and performance. See Table 2.
For the detailed design of the SwinV2 block, we refer to the Swin-Transformer-V2 paper [30]. For AFNO-Swin Fusion block, we follow similar designs from Cross-ViT [31], with the cross section replaced by a FFT truncation operation.
@@ B.1 Model Design
where $K$ is the number of initial conditions, each route has a respective reverse route, making a total of 12 routes. These routes are designed to enable the model to build a effective receptive field (ERF) that can associate information with nearby patches across both spatial and temporal dimensions. We have observed large performance discrepancies when using different groupings of these routes, with the best combination being routes Horizontal-Vertical-Temporal and Temporal-Vertical-Horizontal. Note that while this groupings is the best route grouping based on limited experiments, heuristics and speculative convergence behaviour, a better grouping may still exist.
@@
And found that, the best performance is yielded by model using a full spatial-temporal 3D backbone, we suspect this scan pattern and implementation is insufficient to exploit the spatial relationship but good enough to spatial temporal information.
@@ B.2.1 DEM embedder
We replace the DEM embedder with a all-linear module, slight degradation of performance on 2t variables higher than 150 hPa is observed, see Figure 6
@@ B.2.3 Mamba3D
We replace spatial-temporal Mamba3D blocks with vanilla Mamba2D, which comes natively with VMamba, and observed a slight performance degradation across all variables, with geopotential z and variables above 150 hPa seeing the most improvements, see Figure 8
@@ C.3 Spatial Distribution of Error
In the illustrated examples, we observed degraded performance along the edges and in long lead times, suggesting the LBCs still played a role in limiting the performance of LAM. We also note the performance improvements on mountainous areas in a majority of variables, with the exception of u@850 and v@850. Large error differential spikes can also be found in north-western China, where the climate tends to be drier, this could be a manifestation of the overestimation of water contents in the mamba model.
@@ C.5 Spectra
We follow WeatherBench2[34] and adopt its zonal energy spectrum calculation in a limited area setting. See power spectrum analysis of 8 headline variables in Figure 37. The DLWP-LAMs have markedly better finer detail retention than its host model FourCastNet(SFNO), with MetMamba taking first place among the majority of evaluated variables.

### v2

@@ 1 Introduction
Trained on European Centre for Medium-range Weather Forecast (ECMWF)’s reanalysis product ECMWF Reanalysis v5 (ERA5) dataset, deep learning based weather prediction (DLWP) models have shown promising performance, with FourCastNet[5] being the first data-driven model that directly competes with ECMWF’s Integrate Forecasting System (IFS), followed by a series of other models[6][7][8][9] that outperforms IFS in various metrics [10], these models demonstrates excellent ability in capturing global weather trends, with a fraction of the computation requirement.
Though a coarser grid global forecast is crucial for gaining meaningful insights into longer horizon outlooks, extreme weather events often come in the form of convection, which are sub-grid dynamics not reflected in coarser grid data, even with advancements in fine grid global forecast, such events also requires localized tunning[11], making limited area model (LAM) pivotal in many applications. We conduct this research within the LAM context, partly due to its greatly reduced compute, but also aim to explore the feasibility of deep learning for Limited Area Modeling (LAM).
@@ 2.2 Transformers
Transformers, known for its ability to associate information in various parts of the input data via the attention mechanism[18], are initially designed for natural language processing tasks, have quickly gain wide adoption in computer vision with Vision Transformers[19]. Its quadratic complexity (1) however, poses as a significant hurdle in its application, particularly in vision. Different approaches to alleviate this issue have been proposed, most popular among them is Swin-Transformer[20], which uses a switching window mechanism to reduce the number of required tokens to achieve a global receptive field. Sub-quadratic approximations of attention like Linformer[21] are also a viable surrogate. In the DLWP space, where the compute efficiency is of a lesser concern, the use of transformers is ubiquitous[7][8][22][9][23], favoured by its modeling power. But as noted by [23], the patching step which are used to reduce the number of tokens also leads to degradation in performance.
@@ 2.3 Mamba Model
Mamba[12] model is a class of State Space Models (SSMs), which is originally designed to process continuous signals, bearing close similarity in its formulation to RNNs. (Eq.(2)) Its primary contribution, aside from parallelization and hardware level optimization, is enabling SSMs to process inputs in a time dependent manner, making it a viable alternative to attention [24] Modifying it for vision use cases involves transferring its 1D-Scan pattern to a 2d space[13][25], which achieves better performance than other leading vision transformer models.
@@ 3.1 Dataset
For the lateral boundary condition, we chose the latest iteration of FourCastNet (SFNO), distributed with Earth2Studio[29]. Among models with wide adoption and benchmarks, it is the only one with a continuous 6 hour rollout scheme, specific humidity as its vapor related variable and sufficient years outside of the training set for LBC generation. The LBCs are produced by initializing the model with ERA5 global data from 2018 onward, making a relatively smaller (2018~2021) for LBCs adaptation training, avoiding the years where the global model was trained on. See Appendix A for details regarding the dataset used.
Figure 1: (a) general model architecture: the DLWP-LAM takes primary inputs: two initial conditions $IC_{0}$ at initialization time $t_{0}$ and $IC_{-1}$ at previous (6 hours before) initialization time $t_{0-6h}$, and auxiliary inputs: constants from Digital Elevation Model (DEM) and elapsed time, they are processed by a DEM specific embedder, a conditioning module or by simple concatenation. The input tensor is then processed by $N$ blocks. The output is decoded with a simple linear layer and pixel shuffle operation. (b) DEM Embedder: an embedder that processes different information from the DEM accordingly. (c) AdaLN Block: a block that integrates elapsed year time via the use of adaLN, this will regulate all blocks in the process step.
@@ 3.2 Models
This is a Swin-Transformer-V2 block used in Swin-Transformer-V2[30] and also by FuXi[8], with no modifications, we refer to their paper for implementation details. LAM using this block is referred to as MetSwin.
@@
We design a dual branch model that mixes latent information between Fourier Neural Operator and SwinV2 branches using a FFT cross attention mechanism [31], achieving better results than using either of the single branch models, while alleviating the limitation of AFNO in the limited area setting. LAM using this block is referred to as MetSwinAFNOFusion.
@@
We adopt the cross-scan and cross-merge methods used by VMamba[13], and extend these methods to the 3 dimensional space, to better integrate information from spatial and temporal dimensions, similarly, a 3D depth-wise convolution module is also used. We delineate the design choices of Mamba-3D in Appendix B.2.3 LAM using this block is referred to as MetMamba or MetMamba-3D, its 2 dimensional counterpart with basic 2D spatial scan pattern is referred to as MetMamba-2D.
@@ 3.6 Evaluation
We evaluate all LAM models, FourCastNet(SFNO) and relevant ablation studies with initial conditions from test year 2021. In the total 1439 initial conditions, only 1418 of them are used, with the first 1 and last 20 initial conditions reserved for model initialization and forecast truncation. The models are run and evaluated for 20 steps. Only Root Mean Squared Error (RMSE) is provided as a performance metric for simplicity. We note that this metric is calculated following the convention by WeatherBench[34], with averaging as the last step of the calculation, latitude weighting scheme is not used.
@@ 3.7 Environment
These models are trained on two Nvidia A100-40GB cards utilizing PyTorch’s Fully Sharded Distributed Parallel, gradient checkpointing and Automatic Mixed Precision in bf16 datatype.
@@ 5 Discussion
The Mamba block, when used as a drop-in replacement in our existing LAMs offers a notable performance uplift. This can be attributed to its ability to both generate global receptive field and associate nearby tokens more efficiently than convolution networks. A reduced direct association between disjointed and distant tokens than vision transformers, which we suspect is less common in the field of meteorology, could also be a contributing factor. Its mathematical formulation being more suited to process continuous signals is another valid hypothesis. We also note that the cross scan pattern for VMamba and MetMamba is natively transferable and more suitable to a global geometry.
@@
: Previous models have used simple linear layers to process elevation, soil, land-sea mask and other time related information. We use specialized modules according to process each type of data, See Appendix B.2 for effects of these modules.
@@ Appendix A Data Details
Digital Elevation Model data is sampled from the Weather Research and Forecast (WRF) model, containing land-sea mask, soil type, snow albedo etc. Positional information is added by calculating the sine and cosine normalized latitude and longitude. (See Eq.(A) and Eq.(6)) Time of day is added as a group of 2 dimensional sine and cosine normalized local time. Time of year is added as a group of 1 dimensional sine and cosine normalized current UTC time (in seconds) divided by total seconds per year.
@@ Appendix B Model Details
We delineate the design and implementation details of the our MetMamba model in this section. The main difference to two other types of DLWP-LAMs is MetMamba’s ability to process spatial-temporal inputs natively, without the need for an additional 3D convolution encoder. Aside from sharing the same structure, auxiliary information encoder and decoder, all DLWP-LAMs share the same embedding size and similar parameter count. The SwinV2 based LAM serves as a baseline, the AFNO-Swin based LAM serves as a incremental update with relatively higher efficiency and performance. See Table 2.
For the detailed design of the SwinV2 block, we refer to the Swin-Transformer-V2 paper [30]. For AFNO-Swin Fusion block, we follow the dual branch design from Cross-ViT [31], with one AFNO block for each of the four model stages, the branches are mixed together with FFT truncation operations.
@@ B.1 Model Design
where $K$ is the number of initial conditions, each route has a respective reverse route, making a total of 12 routes. These routes are designed to enable the model to build a effective receptive field (ERF) that can associate information with nearby patches across both spatial and temporal dimensions. We have observed large performance discrepancies when using different groupings of these routes, with the best combination being routes Horizontal-Vertical-Temporal and Temporal-Vertical-Horizontal. Note that while this grouping is the best route grouping based on limited experiments, heuristics and speculative convergence behaviour, a better grouping may still exist.
@@
And we have found that, the best performance is yielded by model using a full spatial-temporal 3D backbone, we suspect this scan pattern and implementation is insufficient to exploit the spatial relationship but good enough to spatial temporal information.
@@ B.2.1 DEM embedder
We replace the DEM embedder with a all-linear module, slight degradation of performance on 2t and variables higher than 150 hPa is observed, see Figure 6
@@ B.2.3 Mamba3D
We replace spatial-temporal Mamba3D blocks with vanilla Mamba2D, which comes natively with VMamba, and observed a slight performance degradation across all variables, with geopotential z and variables above 150 hPa seeing the most regression, see Figure 8
@@ C.3 Spatial Distribution of Error
In the illustrated examples, we observed degraded performance along the edges and in long lead times, suggesting the LBCs still played a role in limiting the performance of LAM. We also note the performance improvements on mountainous areas in a majority of variables, with the exception of u@850 and v@850. Large error differential spikes can also be found in north-western China, where the climate tends to be drier, this could be a manifestation of the overestimation of water contents in the Mamba based model.
@@ C.5 Spectra
We follow WeatherBench2[34] and adopt its zonal energy spectrum calculation in a limited area setting. See power spectrum analysis of 8 headline variables in Figure 37. The DLWP-LAMs have markedly better finer detail retention than its host model FourCastNet(SFNO), with MetMamba taking first place among majority of evaluated variables.